OpenStudy (anonymous):

determine the best method to solve each system of equation.then solve the system. 1.x+2y=3; x+y=1 2.m+6n=-8; m=2n+8 3.a-b=6; a=2b+7 4.4x+y=15; -x-3x=-12 5.3c-d=14; c-d=2 6.x+2y=-9; y=4x

OpenStudy (anonymous):

easy method is to substitute the value of one variable of one equation in the other equation,

OpenStudy (anonymous):

suppose for 1 x=3-2y substitute x=3-2y in the second equation x+y=1

OpenStudy (anonymous):

which gives (3-2y) + y = 1 \[\rightarrow\] 3 - y = 1 y = 2 now from my previous comment x = 3 - 2y = 3 - 2(2) = -1

OpenStudy (anonymous):

thanks